# Learning to Learn with Contrastive Meta-Objective

### v3

@@ Appendix A Complexity Analysis
A.1 In-Context Learning
For ICL model like LLM, ConML does not obtain model representation by explicit model parameters, but by simply adding an additional token to the forward-pass ($u$ in eqn(4)). Which means pre-training a LLM with ConML only reuqires $K/n$ ($K$: subset sampling number, $n$: averge sentence length, typically $K/n<<1$) times computation comparing with pre-training a LLM without ConML, regardless of the model size.
@@ A.2 Typical Meta-Learners
To summarize, for ICL model like LLMs, the complexity to pre-train with ConML is $\frac{n+K}{n}\approx 1$ times the complexity to pre-train without ConML. For typical meta-learners, the additional introduced objective in ConML is comparably less than the complexity of the original meta-training objective, which verifies the empirical computation cost presented in Table 4 in main text.
@@ Appendix B Provable Benefits for Generalization
Here we provide the proof of the following lemma and theorem.
Lemma 3.
@@
Theorem 2.
@@
(iv) Consider $g(;\theta^{*}_{\mathcal{L}_{c}})$, where $\mathcal{L}_{c}$ is defined as $d^{in}-d^{out}$ and under cosine distance, as we did in all experiments in Section 4. Easy to prove that with expressive enough $g$, $\mathcal{L}_{c}$ is minimized only if $P_{g(\theta)}$ is symmetric, i.e., $\forall X\in R^{d},P_{g(\theta^{*}_{\mathcal{L}_{c}})}(X)=P_{g(\theta^{*}_{\mathcal{L}_{c}})}(-X)$.
@@ Appendix C Specifications of Meta-Learning with ConML
Algorithm 5 In-Context Learning with ConML (ICL w/ ConML).
@@
Appendix D In-Context Learning with ConML
D.1 In-Context Learning
In-context learning (ICL) is first proposed for large language models [8], where examples in a task are integrated into the prompt (input-output pairs) and given a new query input, the language model can generate the corresponding output. This approach allows pre-trained model to address new tasks without fine-tuning the model. For example, given "happy->positive; sad->negative; blue->", the model can output "negative", while given "green->cool; yellow->warm; blue->" the model can output "cool". ICL has the ability to learn from the prompt. Training ICL can be viewed as learning to learn, i.e., meta-learning [37, 20, 31]. More generally, the input and output are not necessarily to be natural language. In ICL, a sequence model $T_{\theta}$ (typically transformer [58]) is trained to map sequence $x_{1},y_{1},x_{2},y_{2},\cdots,x_{m-1},y_{m-1},x_{m}]$ (prompt prefix) to prediction $y_{m}$. Given distribution $P$ of training prompt $t$, then training ICL follows an auto-regressive manner:
@@ D.3 Integrating ConML with ICL
where $u$ is constrained to be in the same shape as $x$, and has consistent value in an episode. The complete algorithm of ConML for ICL is provided in Appendix C.For example, for training a ICL model on linear regression tasks we can choose $u=\bm{1}$, and in pre-training of LLM we can choose u=u="what is this task?". From the perspective of learning to learn, ConML encourages ICL to align and discriminate like it does for conventional meta-learning, while the representations to evaluate inner- and inter- task distance are obtained by probing output rather than explicit model weights. Thus, incorporating ConML into the training process of ICL benefits the fast-adaptation and task-level generalization ability. From the perspective of supervised learning, ConML is performing unsupervised data augmentation that it introduces the dummy input and contrastive objective as additional supervision to train ICL.
@@ Appendix F Experimental Results Obtained Using Different Backbones
Appendix G Experimental Results on Few-shot Molecular Property Prediction
Few-shot molecular property prediction is another important test-bed for meta-learning methods [3, 23, 65, 11, 48]. We use FS-Mol [52], a widely studied benchmark consisting of a large number of diverse tasks. We follow the public data split provided in [52]. Each training set contains 64 labeled molecules and can be imbalanced, where the number of labeled active and inactive molecules may not be equal. The remaining molecules in each task form the validation set.
[…]
We incoporate ConML into state-of-the-art method PACIA, an amortization-based meta-learner, as PACIA w/ ConML. The embedding vectors from $\mathcal{D}^{\text{tr}}$ are input into the hypernetwork, and the output modulates the embedding vectors through FiLM. We set the hyperparameters for ConML as follows: $B=16$, $\phi(a,b)=1-\nicefrac{{a\cdot b}}{{\|a\|\|b\|}}$ (cosine distance), and $\lambda=0.1$. For the sampling strategy $\pi_{\kappa}$ and the number of times $K$, we sample subsets of different sizes for each task, specifically $m\in{4,8,16,32,64}$, with $128/m$ iterations for each size. Each subset contains $m/2$ positive and $m/2$ negative samples, selected randomly. All other hyperparameters related to model structure and training follow the default settings from the benchmark [52].
Table 8 presents the results. PACIA w/ ConML outperforms the state-of-the-art approach across all meta-testing scenarios with different shots. Comparing PACIA w/ ConML to the original PACIA, the impact of ConML is notably significant.
[…]
2-shot 4-shot 8-shot 16-shot
MAML $009\pm{.006} .$$125\pm{.009} .$$146\pm{.007} .$$159\pm{.009}$
PAR $124\pm{.007} .$$140\pm{.005} .$$149\pm{.009} .$$164\pm{.008}$
ProtoNet $117\pm{.006} .$$142\pm{.007} .$$175\pm{.006} .$$206\pm{.008}$
CNP $139\pm{.004} .$$155\pm{.008} .$$174\pm{.006} .$$187\pm{.009}$
ADKF-IFT $131\pm{.007} .$$166\pm{.005} .$$202\pm{.006} .$${.234}\pm{.009}$
PACIA $142\pm{.007} .$$169\pm{.006} .$$205\pm{.008} .$$236\pm{.008}$
PACIA w/ ConML $\bm{.175}\pm{.006}$ $\bm{.196}\pm{.006}$ $\bm{.218}\pm{.005}$ $\bm{.241}\pm{.007}$
Appendix H Experiment Details
H.1 Model Analysis
@@ H.2 In-Context Learning
Appendix I Related Works
The most closely related works which bring regularization to meta-objective to align and discriminate have been discussed in Section 5. It is demonstrated that they are either problem-specific or learner-specific, while ConML exploits the intrinsic task-identity brought by the mini-batch episodic training, which is problem- and learner-agnostic. The problem-specific works [67, 17, 72, 63, 33] can be incorporated with ConML to bring performance boost as they all work with/under the general mini-batch episodic training, and the learner-specific works [22, 35] can be viewed as specifications of ConML.
Here we introduce about a broader range of related works about learning to learn and contrastive learning.
I.1 Learning to Learn
Meta-learning learns to improve the learning algorithm itself [49], i.e., learns to learn. Popular meta-learning approaches can be roughly divided into three categories [7]: optimization-based, metric-based and amortization-based. Optimization-based approaches [4, 19, 40] focus on learning better optimization strategies for adapting to new tasks. For example MAML [19] learns initial model parameters, where few steps of gradient descent can quickly make adaptaion for specific tasks. Metric-based approaches [59, 51, 53] leverages learned similarity metrics. For example, Prototypical Networks [51] and Matching Networks [59] learn global shared encoders to map training set to embeddings, based on which task-specific model can be built. Amortization-based approaches [21, 45, 6] seek to learn a shared representation across tasks. They amortize the adaptation process by using neural networks to directly infer task-specific parameters from training set. Examples are CNPs [21] and CNAPs [45].
In-context learning (ICL) [8] is designed for large language models, which integrates examples (input-output pairs) in a task and a query input into the prompt, thus the language model can answer the query. Recently, ICL has been studied as a general approach of learning to learn [2, 20, 60, 1], which reduces meta-learning to conventional supervised learning via training a sequence model. It considers training set as context to be provided along with the input to predict, forming a sequence to feed the model. Training such a model can be viewed as an instance of meta-learning [20, 70].
I.2 Contrastive Learning
Contrastive learning is a powerful technique in representation learning [41, 10, 64]. Its primary goal is to learn useful representations, which are invariant to unnecessary details, and preserve as much information as possible. This is achieved by maximizing alignment and discrimination (uniformity) in representation space [64]. In conventional contrastive learning, alignment refers to bringing positive pairs (e.g., augmentations of the same sample [71, 26, 5, 24, 10]) closer together in the learned representation space. By maximizing alignment, the representations are encouraged to be invariant to unneeded noise factors. Discrimination refers to separating negative pairs (e.g., different samples) farther. Maximizing discrimination without any other knowledge results in uniformity, i.e., uniform distribution in the representation space. By maximizing discrimination, the representations are encouraged to preserve as much information of the data as possible [56, 5], benefiting the generalization ability.
While ConML and conventional feature-contrastive learning share the core concept of maximizing alignment and discrimination, ConML extends this idea to the model space within meta-learning and utilizes task identity as additional supervision. This approach enables ConML to address the unique challenges of meta-learning and improve the generalization ability of meta-learners across a wide range of tasks and domains. The differences can be summarized as:
  • Scope of Application: While feature-contrastive learning primarily focuses on unsupervised representation learning in the representation space, ConML extends this concept to the model space within the context of meta-learning. It aims to improve the generalization ability of meta-learners by enhancing their alignment and discrimination capabilities through task-level contrastive learning.
  • Supervision Source: Feature-contrastive learning relies on unlabelled data and their augmentations to learn invariant representations. In contrast, ConML utilizes task identity as an additional supervision signal, allowing it to leverage the inherent structure of the meta-learning setting. This task-level supervision enables ConML to directly optimize the meta-learner’s model space, promoting both alignment within tasks and discrimination between tasks.
  • Objectives: Feature-contrastive learning primarily aims to learn representations that are robust to irrelevant noise and preserve essential information. ConML, while still benefiting from these properties, focuses on improving the meta-learner’s ability to adapt to new tasks. By optimizing for alignment and discrimination, ConML enhances the meta-learner’s generalization ability, enabling it to quickly learn and adapt to unseen tasks with limited data.

### v4

@@ Appendix A Complexity Analysis
A.1 ICL
For ICL model like LLM, ConML does not obtain model representation by explicit model parameters, but by simply adding an additional token to the forward-pass ($u$ in (4)). Which means pretraining a LLM with ConML only requires $K/n$ ($K$: subset sampling number, $n$: average sentence length, typically $K/n<<1$) times computation comparing with pretraining a LLM without ConML, regardless of the model size.
@@ A.2 Typical Meta-Learners
To summarize, for ICL model like LLMs, the complexity to pretrain with ConML is $\frac{n+K}{n}\approx 1$ times the complexity to pretrain without ConML. For typical meta-learners, the additional introduced objective in ConML is comparably less than the complexity of the original meta-training objective, which verifies the empirical computation cost presented in Table 4 in main text.
@@ Appendix B Provable Benefits for Generalization
Here we first provide the proof of Lemma 1 which shows $U_{p(\tau)}(\theta)$ is an upper bound of the excess risk of meta-learning $\Delta\epsilon_{p(\tau)}(\theta)$, and then the proof of Theorem 1 which shows minimizing contrastive meta-objective is minimizing $U_{p(\tau)}(\theta)$.
We need two preliminary results:
Lemma 2 (Upper Bound from [28]).
$\forall\theta$, $\Delta\epsilon_{p(\tau)}(\theta)\leq U_{p(\tau)}(\theta)$, where
@@
with $C_{1},C_{2}>0,\frac{dC_{1}}{dg}=\frac{dC_{2}}{dg}=0$ .
Lemma 3 (Universal Approximation of MLP from [21]).
Let $\sigma:\mathbb{R}\to\mathbb{R}$ be a non-constant, bounded, and continuous function. Let $K$ be a compact subset of $\mathbb{R}^{n}$. The set of real-valued continuous functions on $K$ is denoted by $C(K)$. For any function $f\in C(K)$ and for any error tolerance $\delta>0$, there exists an integer $N$ (the number of neurons in the hidden layer), and real constants $v_{i},b_{i}\in\mathbb{R}$ and vectors $w_{i}\in\mathbb{R}^{n}$ for $i=1,\dots,N$, such that we can define the MLP output function $F:K\to\mathbb{R}$ as:
$F(x)=\sum_{i=1}^{N}v_{i}\sigma(w_{i}^{T}x+b_{i})$
which satisfies $\forall x\in K$ and $|f(x)-F(x)|<\delta$.
B.1 Proof of Lemma 1
@@
We use the Universal Approximation of MLP to bound the $\psi$ in $\Phi_{p(\tau)}(\theta)$.
For the model space $\mathcal{H}$ (recall $h=g(\mathcal{D};\theta)\in\mathcal{H}$) is closed and bounded by the meta-learner’s hypothesis, which is a compact set on $\mathbb{R}^{|\mathcal{H}|}$, we can apply Theorem 3 on $\psi$ in $\Phi_{p(\tau)}(\theta)$. There exists two-layer MLP $F$ with a nonlinear activation function $\sigma$. Then
$\forall X_{g}\in\mathcal{H},\quad|\psi(X_{g})-F(X_{g})|<\delta,$ (5)
where $X_{g}$ represents the parameter vector learned by $g(;\theta)$ from sample $(x,y)$.
[…]
As $\delta$ can be arbitrarily small by selecting large enough $N$, we approximately rewrite (5) as $\psi(X_{g})=F(X_{g})$ which is not completely rigorous but no harm to our proof. We have
[…]
where $P_{g(\theta)(\tau,x,y)}$ is the distribution of $X_{g}$ output by meta-learner $g(;\theta)$ on defined $p(\tau),x,y$.
Let the activation function $\sigma$ has Lipschitz constant $L_{\sigma}$ and $\sigma(0)=0$. We have
[…]
So we have
[…]
So $U_{p(\tau)}(\theta)$ is a upper bound of he the excess risk of a meta-learner $\Delta\epsilon_{p(\tau)}(\theta)$. ∎
B.2 Proof of Theorem 1
Proof.
On the one hand, we have
[…]
$\geq C_{2},$
where the minimum of $U_{p(\tau)}(\theta)$
$U^{*}_{p(\tau)}(\theta)=C_{2}$
is achieved if and only if
[…]
which is a achieved if and only if
$\forall X\in\mathcal{H},P_{g(\theta)(\tau,x,y)}(X)=P_{g(\theta)(\tau,x,y)}(-X),$
i.e., $P_{g(\theta)(\tau,x,y)}$ is symmetric.
On the other hand, by definition we have
$\mathcal{L}_{c}=$ $d^{in}-d^{out}$
@@ B.2 Proof of Theorem 1
For arbitrary subset sampling strategy $\pi_{\kappa}$, we have
$\min_{\theta}\mathcal{L}_{c}\geq-1,$
where the minimum of $\mathcal{L}_{c}$
$\mathcal{L}^{*}_{c}=-1,$
is achieved if and only if
@@
[…]
which is achieved if and only if
$\forall\tau,P_{g(\theta)}(\tau)=P_{g(\theta)}(-\tau),P_{g(\theta,\pi_{\kappa})(x,y,\kappa\mid\tau)}(X\mid\tau)=\delta_{\tau}(X),$
where $\delta_{\tau}$ is the Dirac-delta function centered on $\tau$. Then
$P_{g(\theta,\pi_{\kappa})(\tau,x,y,\kappa)}(X)=$ $P_{g(\theta)}(\tau)P_{g(\theta,\pi_{\kappa})(x,y,\kappa\mid\tau)}(X\mid\tau)$
$=$ $P_{g(\theta)}(\tau)\delta_{\tau}(X)=P_{g(\theta)}(\tau)\delta_{\tau}(-X)$
$=$ $P_{g(\theta,\pi_{\kappa})(\tau,x,y,\kappa)}(-X).$
Combining both hands, we have
$\theta^{*}_{\mathcal{L}_{c}}=\arg\min_{\theta}\mathcal{L}_{c}(g(;\theta),p(\tau))$
$\Rightarrow P_{g(\theta^{*}_{\mathcal{L}_{c}},\pi_{\kappa})(\tau,x,y,\kappa)}(X)=P_{g(\theta^{*}_{\mathcal{L}_{c}},\pi_{\kappa})(\tau,x,y,\kappa)}(-X)$
$\Leftrightarrow U_{p(\tau)}(\theta^{*}_{\mathcal{L}_{c}})=C_{2}=\min_{\theta}U_{p(\tau)}(\theta)$
@@ Appendix C Specifications of Meta-Learning with ConML
Algorithm 5 ICL with ConML (ICL w/ ConML).
@@
Appendix D ICL with ConML
D.1 ICL
ICL is first proposed for LLMs [6], where examples in a task are integrated into the prompt (input-output pairs) and given a new query input, the language model can generate the corresponding output. This approach allows pretrained model to address new tasks without fine-tuning the model. For example, given "happy->positive; sad->negative; blue->", the model can output "negative", while given "green->cool; yellow->warm; blue->" the model can output "cool". ICL has the ability to learn from the prompt. Training ICL can be viewed as learning to learn, i.e., meta-learning [29, 16, 24]. More generally, the input and output are not necessarily to be natural language. In ICL, a sequence model $T_{\theta}$ (typically transformer [45]) is trained to map sequence $x_{1},y_{1},x_{2},y_{2},\cdots,x_{m-1},y_{m-1},x_{m}]$ (prompt prefix) to prediction $y_{m}$. Given distribution $P$ of training prompt $t$, then training ICL follows an auto-regressive manner:
@@ D.3 Integrating ConML with ICL
where $u$ is constrained to be in the same shape as $x$, and has consistent value in an episode. The complete algorithm of ConML for ICL is in Appendix C.For example, for training a ICL model on linear regression tasks we can choose $u=\bm{1}$, and in pretraining of LLM we can choose u=u="what is this task?". From the perspective of learning to learn, ConML encourages ICL to align and discriminate like it does for conventional meta-learning, while the representations to evaluate inner- and inter- task distance are obtained by probing output rather than explicit model weights. Thus, incorporating ConML into the training process of ICL benefits the fast-adaptation and task-level generalization ability. From the perspective of supervised learning, ConML is performing unsupervised data augmentation that it introduces the dummy input and contrastive objective as additional supervision to train ICL.
@@ Appendix F Experimental Results Obtained Using Different Backbones
Appendix G Implementation Details
G.1 Model Analysis